A minor league baseball team raised the average price of its tickets from S8 to $8.TO and found that average attendance at its games dropped from 5,200 to 4,800. Using the arc elasticity of demand formula, the price elasticity of demand for tickets is(Express your answer as a real number rounded to two decimal places. Dont forget the negative sign)
Answer #1

Ans: -0.95

Explanation:

Here, P1 = 8, Q1 = 5200,

P2 = 8.70, Q2 = 4800

PED = (Q2 - Q1) / (P2 - P1) * (P1 + P2) / (Q1 + Q2)

= (4800 - 5200) / (8.70 - 8) * (8 + 8.70) / (4800 + 5200)

= (-400 / 0.70) * (16.70 / 10000)

= -6680 / 7000

= -0.95